(9 - 5)2 + 5 + 4² =
Answer by atif.muhammad(135)

(9 - 5)2 + 5 + 4² =
I think you meant
(9 - 5)² + 5 + 4² = 4² + 5 + 4² = 16 + 5 + 16 = 21 + 16 = 37
If you didn't mean this then:
(9 - 5)2 + 5 + 4² = 4(2) + 5 + 4² = 8 + 5 + 16 = 13 + 16 = 29
Whichever one is the correct one.